If the price of electrical energy is $0.14 per kilowatt·hour, what is the cost of using electrical energy to heat the water in a swimming pool (12.0 m x 9.00 m x 1.50 m) from 13 to 28 °C?
Answer #1

The mass of water = 162,000 kg (162 m3 x 1000 kg)

The specific heat for water is 4.18J/gC or 4.18 kJ/kgC
energy = 162,000 kg * 4.18 kJ/kgC * 15 deg C = 1.016 * 10^7 kJ

Now 1 kWh * 1 (kJ/sec)/kW * 3600 sec/h = 3600 kJ
So 1 kWh = 3600 kJ and costs you 14 cents.

1.016 * 10^7 kJ / 3600 kJ * $0.14 = $395.11